Sir,
10 years ago a man has 6 times as old as his daughter. After 10 years he will be twise as old as his daughter. Find their present age.
thank you-Bindu

Answer by swincher4391(1107)

let m be the age of the man let d be the age of the daughter
(m-10) = 6*(d-10)
m+10 = 2*(d+10)

m-10 = 6d - 60
m+10 = 2d + 20
m = 6d -50 (solving first one for m)

6d-50 + 10 = 2d + 20
6d -40 = 2d + 20
4d = 60
d = 15
m = 6(15) - 50 = 40
The man is 40, his daughter is 15.